# Bridging Classical and Quantum Machine Learning: Knowledge Transfer from Classical to Quantum Neural Networks Using Knowledge Distillation

### v1

@@ Abstract
Very recently, studies have shown that quantum neural networks surpass classical neural networks in tasks like image classification when a similar number of learnable parameters are used. However, the development and optimization of quantum models are currently hindered by issues such as qubit instability and limited qubit availability, leading to error-prone systems with weak performance. On the other hand, classical models can exhibit high-performance owing to substantial resource availability. As a result, more studies have been focusing on hybrid classical-quantum integration to leverage the advantages of both paradigms. A line of research particularly focuses on transfer learning through classical-quantum integration or quantum-quantum approaches. Unlike previous studies, this paper introduces a new method to transfer knowledge from classical to quantum neural networks using knowledge distillation, effectively bridging the gap between classical machine learning and emergent quantum computing techniques. We adapt classical convolutional neural network (CNN) architectures like LeNet and AlexNet to serve as teacher networks, facilitating the training of student quantum models by sending supervisory signals during backpropagation through KL-divergence. The approach yields significant performance improvements for the quantum models by solely depending on classical CNNs, with quantum models achieving an average accuracy improvement of 0.80% on the MNIST dataset and 5.40% on the more complex FashionMNIST dataset. Applying this technique eliminates the cumbersome training of huge quantum models for transfer learning in resource-constrained settings and enables re-using existing pre-trained classical models to improve performance. Thus, this study paves the way for future research in quantum machine learning (QML) by positioning knowledge distillation as a core technique for advancing QML applications.
Keywords Parameterized quantum circuits $\cdot$ Quantum superposition $\cdot$ Knowledge distillation $\cdot$ Convolutional neural networks  $\cdot$ Classification
@@ 1 Introduction
footnotetext: * Corresponding author. E-mail address: mahdy.chowdhury@northsouth.edu (M.R.C. Mahdy).
Quantum machine learning (QML) holds great promise to revolutionize computational paradigms by offering unprecedented computational speedups for certain tasks through parallel computation leveraging quantum superposition and entanglement. In particular, quantum neural networks (QNNs) have shown promising results compared to classical approaches by outperforming classical models in image classification tasks and converging faster when similar number of learnable parameters are used [1, 2]. Moreover, well-constructed QNNs can have a substantially higher effective dimension than classical neural networks, implying a greater capacity for quantum networks to handle complex data structures and tasks [3]. Despite the promising advancements recently, the infancy of QML poses intrinsic challenges. Although they are theoretically potent, the development and optimization of QNNs are often constrained by factors such as quantum bit (qubit) stability, coherence times, and restricted qubit availability [4, 5]. This results in a system that is error-prone, limiting the complexity and expressive power of models that can be implemented and evaluated on actual quantum hardware [6, 7, 8]. This embryonic stage of QML unavoidably collides with the immense computational burden required to train highly-accurate models, a situation further worsened by the inherently stochastic nature of quantum computations and the concurrent need for substantial resources to manage and mitigate resultant quantum noise.
Existing approaches to handle the errors and resource-limitations of QNNs include neural error mitigation, variational quantum-neural hybrid error mitigation, resource-efficient quantum circuit design, quantum-to-quantum transfer learning and classical-to-quantum transfer learning. However, each of these approaches has its own limitations. Neural error mitigation requires a large amount of training data and computational resources to train the neural networks that correct the quantum errors[9]. Variational quantum-neural hybrid error mitigation introduces additional variational parameters that need to be optimized, which can increase the complexity and instability of the quantum optimization process. Resource-efficient quantum circuit design depends on the specific choice of the linear transformation and the quantum feature map, and may not be applicable to arbitrary quantum tasks [10]. Quantum-to-quantum transfer learning assumes that the source and target quantum systems or tasks are sufficiently similar, and may not work well for heterogeneous or diverse quantum domains. Classical-to-quantum transfer learning relies on the classical neural network to extract meaningful features from the data, which may not capture the quantum correlations or entanglement that are essential for quantum advantages.
Transfer learning, especially using knowledge distillation [11], effectively navigates these challenges by offering a straightforward alternative where large pre-trained models share their learned knowledge with smaller ones in a teacher-student setup. This approach obviates the necessity of burdensome training and error mitigation to improve model performance by depending solely on the logits of large pre-trained models, assuaging the computational cost. Furthermore, this method also circumvents the limitations of the available resources by enabling smaller models to inherit robust, pre-optimized decision boundaries and representational features from the teacher models. KD can be used to transfer the knowledge learned by a large or complex neural network (the teacher) to a smaller or simpler neural network (the student). It can also reduce the noise sensitivity of the student network, and enable cross-domain or cross-hardware applications. In addition, it can enhance the interpretability and explainability of the student neural network, by revealing the underlying physical rules or symmetries that govern the performance enhancement. As a result, knowledge distillation can be used and applied as a promising technique for neural network design and optimization.
Knowledge distillation in neural networks can be accomplished in various ways: (1) Response-based knowledge distillation [12]: This approach transfers the knowledge from the teacher network to the student network by matching their outputs or responses on a given dataset. The student network learns to mimic the teacher network’s predictions, probabilities, or logits, and thus inherits the teacher network’s generalization ability. (2) Feature-based knowledge distillation [12]: This approach transfers the knowledge from the teacher network to the student network by matching their intermediate features or representations. The student network learns to extract similar features as the teacher network, and thus captures the teacher network’s discriminative power. (3) Relation-based knowledge distillation [12]: This approach transfers the knowledge from the teacher network to the student network by matching their relations or interactions among the inputs, outputs, or features. The student network learns to preserve the same relations as the teacher network, and thus acquires the teacher network’s structural or semantic information.
Recent studies have explored transfer learning in quantum neural networks with respoonse-based knowledge distillation using a Quantum-to-quantum approach [9]. A large domain of Knowledge Distillation remains unexplored: Classical-to-quantum knowledge transfer. Given the constraints of quantum machine learning, it is crucial to explore the possibilities that a quantum model can enhance its performance solely based on the outputs of classical neural networks. This leads us to the research questions of this study, based on which we design the methodology, experiment rigorously, analyze and conclude our findings in this paper:
@@
  • Research Question 1: Can we transfer the knowledge from a classical neural network (the teacher) to a quantum neural network (the student) using knowledge distillation?
  • Research Question 2: How does knowledge distillation from classical neural network to quantum neural network affect the performance of the quantum model?
  • Research Question 3: Is it possible to improve the performance of the quantum models upon knowledge distillation from classical neural networks?
In this paper, we show theoretically and experimentally that knowledge transfer from classical to quantum models using knowledge distillation indeed improves the performance of the quantum models drastically. This method obviates the dependency of QNNs upon larger QNNs, thus eliminating the necessity of hectic model training, quantum error-mitigation and resource handling for improving the performance of quantum models. To demonstrate our claim, we conduct extensive experiments on two widely used datasets: MNIST [13] and FashionMNIST [14]. We use multiple classical convolutional neural networks as teacher models to train on these datasets and generate the outputs for use by the quantum student models. We also use multiple quantum student models like 3-qubit and 4-qubit variational quantum circuits, to learn from the outputs of the classical teacher models and improve performance. We compare the performance of the quantum student models with and without knowledge distillation, and show that knowledge distillation can significantly improve the accuracy, robustness, and efficiency of the quantum models. Moreover, we also use ensembles of the classical teacher models, and investigate how the performance of the quantum student models varies with the number of teacher models used. We find that using more teacher models can provide more diverse and complementary information for the quantum student models, and thus further enhance the knowledge transfer effect.
The methodology followed in this study is visualized in  figure 1. The figure shows that after following data pre-processing steps for classical and quantum models respectively, the classical data is passed through a classical network containing teacher CNN model(s), whereas the quantum data is passed through student QNN model(s). After that, using the outputs of the models, the KL-Divergence is calculated and added to the loss of the student model to calculate a total loss which is used to improve performance of the models. Adding the KL-divergence acts as a supervisory signal while updating the parameters of the quantum model. In this teacher-student setup, the student QNN thus tries to mimic the outputs of the teacher CNN and knowledge transfer is accomplished from classical to quantum models solely based on the mimicry.
Figure 1: A visual representation of the research methodology followed in this research. At the very beginning, data pre-processing is carried out individually for classical and quantum model training. Following that, the data is passed through the respective models. A pre-trained CNN is used in the classical network. The outputs from the CNN and the QNN are used to calculate the KL-divergence and add that to the loss of the quantum network to aid in backpropagation.
Contributions. The contributions of this research can be summarized as follows:
  • We propose a novel hybrid knowledge distillation framework to transfer the knowledge from multiple classical teacher models to quantum student models, which establishes the re-usability of existing pre-trained classical models in quantum training and eliminates the necessity of cumbersome QNN training.
  • We systematically investigate the effect of knowledge distillation from classical neural networks to quantum neural networks on the performance of the quantum models.
  • We conduct extensive experiments on two widely used datasets, MNIST and FashionMNIST, and compare the performance of various quantum student models with and without knowledge distillation.
  • We empirically present that knowledge distillation can significantly improve the accuracy, robustness, and efficiency of the quantum models, and also enable cross-domain and cross-hardware applications.
  • We provide insights into the underlying mechanisms and factors that influence the knowledge transfer process, and suggests future directions for research in this area.
The rest of the paper is organized as follows: section 2 discusses the related works, providing context and background for our approach. section 3 describes the methodology, detailing the mathematical foundations and our approach to knowledge distillation. In section 4, we present the experiments and results, along with the datasets used, models utilized and the specifics of our implementation. section 5 presents the findings of this study and dives into their in-depth discussion, interpreting the results and their implications. section 6 concludes the paper and suggests directions for future work, highlighting the potential applications and extensions of our research. Finally, we acknowledge the contributions of the people that have assisted in this research.
@@ 2 Related Works
2.1 Quantum Machine Learning: Problems and Challenges
Quantum Machine Learning (QML) aspires to harness the properties of quantum mechanics to revolutionize computation, but the path is restricted with various challenges and obstacles. The concept of quantum supremacy, where quantum systems perform tasks beyond the reach of classical computers, has seen experimental strides as evidenced by [15] and [16]. However, these pioneering steps also underscore the fragility of quantum states and their vulnerability to decoherence, a problem that Shor’s early work [17] sought to address through quantum error correction, a technique still pivotal in contemporary research [18]. The transient stability of qubits, a topic explored by [19], and the intricate challenge of manipulating quantum states, as demonstrated by [20], add layers of complexity to quantum computing.
@@ 2.1 Quantum Machine Learning: Problems and Challenges
2.2 Error Mitigation and Resource Handling in Quantum Computing
@@
The pursuit of reliable quantum computing is marked by a need to address the fundamental limits of quantum error mitigation. Works such as those by [28] provide a foundational understanding of the theoretical bounds in quantum systems, crucial for error correction techniques in neural networks. This is further explored through the lens of scalability and statistics of errors in recent advancements within noisy intermediate-scale quantum (NISQ) technologies [29]. Reference-state error mitigation strategies [30] have shown promise in improving the accuracy of quantum computations, which is vital for the fidelity of knowledge distillation processes. Moreover, the extension of computational reach through error mitigation in noisy quantum processors [31] has direct implications for the robust execution of quantum neural networks. The integration of learning-based techniques for quantum error mitigation [32] parallels the adaptive nature of knowledge distillation, suggesting a methodological symmetry that could be harnessed for transferring knowledge. The evaluation and benchmarking of error mitigation techniques are critical for determining the most effective methods for application in quantum neural networks [33]. This selection is further informed by the development of a unified approach to data-driven error mitigation [34], offering a pathway to harnessing empirical data in the error correction process. Techniques for mitigating errors in quantum approximate optimization [35] can also inform the optimization strategies within the knowledge distillation framework.
@@
On the front of resource handling, the importance of optimizing resource efficiencies is underscored by [36], drawing attention to the balance between computational capabilities and resource constraints. This optimization is directly applicable to the development of quantum neural networks, where the allocation and management of quantum resources are crucial. Furthermore, the critical challenges of scaling quantum computation, as outlined by [37], must be addressed to enable the practical implementation of fault-tolerant quantum neural networks. Insights into the cross-discipline theoretical research on quantum computing, such as those presented by [38], highlight the urgency and significance of resource management in the context of knowledge distillation. Additionally, the potential applications of quantum computing in specific domains like renewable energy optimization [39] illustrate the diverse utility and the resource handling strategies that may be required for energy-efficient quantum neural network operations. Differently from previous approaches, we completely nullify the need for QNN optimization using error mitigation or resource handling. Rather, keeping the student quantum model small and static, we depend fully on pre-trained classical models and their ensembles to improve the performance of the QNN, thus addressing the challenges of QML in a new, different way.
2.3 Transfer Learning in Quantum Machine Learning
@@
Transfer learning has emerged as a potent strategy in classical machine learning, enabling models trained on one task to be repurposed for another related task [40, 41]. In quantum machine learning (QML), this approach promises to address scalability and adaptability challenges that are crucial in the practical application of neural-network quantum states (NNQS) [40]. Scalability, a primary concern in QML, is particularly relevant for tasks that require models to extrapolate learned behaviors to larger quantum systems. Protocols inspired by physics have shown that features learned from smaller systems can significantly accelerate and refine the learning process when applied to more complex systems, as evidenced in the one-dimensional transverse field Ising and Heisenberg XXZ models [40]. The versatility of transfer learning is further showcased by its application to the optimization of neural network potentials, exemplified by ANI-1x, which was retrained to near gold-standard quantum mechanical accuracy. This highlights transfer learning’s broad potential across disciplines, including materials science, biology, and chemistry [41, 42].
@@
Integrating classical and quantum neural networks has extended transfer learning protocols into classical-quantum hybrid domains. By combining pre-trained classical networks with variational quantum circuits, researchers have opened up new possibilities for processing high-dimensional data such as images before entangling them within the quantum processing framework [43]. This hybrid approach not only utilizes the strengths of classical networks in data handling but also harnesses the quantum advantages for feature processing, potentially revolutionizing fields like image recognition and quantum state classification [43, 44]. Moreover, the concept of knowledge distillation has been introduced into the realm of QML, allowing the creation of new quantum neural networks (QNNs) through approximate synthesis [9]. This methodology enables the reduction of circuit layers or the alteration of gate sets without the necessity of training from scratch. Empirical analyses suggest that significant reductions in circuit complexity can be achieved while simultaneously improving accuracy, even under noisy conditions, underscoring the effectiveness of transfer learning techniques in the development of QNNs [9].
@@
Quantum Convolutional Neural Networks (QCNNs) are gaining popularity in both quantum and classical data classification recently [50]. The adoption of transfer learning in QML is proposed to circumvent the scalability constraints of quantum circuits in the noisy intermediate-scale quantum (NISQ) era. For instance, research has shown that small QCNNs can effectively solve complex classification tasks by leveraging pre-trained classical CNNs, thus bypassing the need for expansive quantum circuitry [49]. This approach is validated through numerical simulations of QCNNs, which demonstrate a superior classification accuracy over classical models when pre-trained on different datasets. The synergy between classical and quantum networks is evident when a classical CNN, trained on FashionMNIST data, is utilized to enhance the performance of a QCNN for MNIST data classification [50]. The empirical results advocate for transfer learning from classical to quantum CNNs, revealing a performance that significantly outshines classical transfer learning models under analogous training conditions. These findings not only elucidate the potential of knowledge transfer to optimize the utilization of QCNNs in the NISQ era but also underscore the practicality of hybrid classical-quantum models in overcoming the limitations posed by the present-day quantum technologies. Driven by these findings, our approach is designed to transfer knowledge from classical CNNs to quantum models. However, differently from [49] and [50], we use knowledge distillation and show empirically that knowledge transfer is possible across homogeneous datasets in quantum models by simply trying to mimic the soft targets of classical CNNs instead of hectic training by freezing the pre-trained model and transferring the knowledge to a QNN during training in the previous approaches.
3 Methodology
3.1 Problem Formulation and Integration of Knowledge Distillation
Given a dataset $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$ where $\mathbf{x}_{i}\in\mathcal{X}$ and $y_{i}\in\mathcal{Y}$, the task is to train the student QNN to approximate the mapping of the teacher model. The performance of the QNN can be evaluated by a loss function $\mathcal{L}:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$, which measures the discrepancy between the predictions of the QNN and the ground truth labels. The objective function for the QNN is given in  equation 1 as:
The loss of the model outputs, $\mathcal{L}$, is calculated using cross-entropy loss [51] according to  equation 2:
Knowledge Distillation (KD) aims to transfer knowledge from the teacher model to the student model by minimizing a distillation loss $\mathcal{L}_{KD}$. The KD process involves a soft target $\tau$ which controls the smoothness of the teacher’s output probability distribution. The distillation loss can be defined using the Kullback-Leibler (KL) divergence [52] in  equation 3 as:
where $\sigma$ denotes the softmax function and KL denotes the Kullback-Leibler divergence. The softmax function is given according to  equation 4 by:
$\sigma(\mathbf{z})_{i}=\frac{e^{z_{i}}}{\sum_{j=1}^{K}e^{z_{j}}}$ (4)
And, the KL-divergence is calculated as per  equation 5:
$KL(P||Q)=\sum_{i=1}^{K}P(i)\log\left(\frac{P(i)}{Q(i)}\right)$ (5)
where $P$ and $Q$ are discrete probability distributions. $P(i)$ represents the probability of the $i$-th class according to the true label distribution, typically obtained from the softmax function of the teacher model $\sigma(f_{t}(\mathbf{x}_{i}))$, and $Q(i)$ is the probability of the $i$-th class according to the predicted label distribution from the student model $\sigma(f_{s}(\mathbf{x}_{i}))$. $K$ is the total number of classes. The KL divergence measures the difference between two probability distributions over the same variable. The total loss for the student QNN, incorporating both the original objective and the distillation loss, is given in  equation 6 by:
$\mathcal{O}_{total}(f_{s})=\mathcal{O}(f_{s})+\lambda\mathcal{L}_{KD}(f_{s},f_{t}),$ (6)
where $\lambda$ is a hyperparameter that balances the importance of the original loss and the distillation loss.
The student QNN is trained by optimizing $\mathcal{O}_{total}$ through a quantum optimization algorithm. The goal of the optimization is to minimize the loss, which can be formulated in  equation 7 as:
$f_{s}^{*}=\arg\min_{f_{s}}\mathcal{O}_{total}(f_{s}).$ (7)

### v2

@@ Abstract
Quantum neural networks (QNNs), harnessing superposition and entanglement, have shown potential to surpass classical methods in complex learning tasks but remain limited by hardware constraints and noisy conditions. In this work, we present a novel framework for transferring knowledge from classical convolutional neural networks (CNNs) to QNNs via knowledge distillation, thereby reducing the need for resource-intensive quantum training and error mitigation. We conduct extensive experiments using two parameterized quantum circuits (PQCs) with 4 and 8 qubits on MNIST, FashionMNIST, and CIFAR10 datasets. The approach demonstrates consistent accuracy improvements attributed to distilled knowledge from larger classical networks. Through ablation studies, we systematically compare the effect of state-of-the-art dimensionality reduction techniques—fully connected layers, center cropping, principal component analysis, and pooling—to compress high-dimensional image data prior to quantum encoding. Our findings reveal that fully connected layers retain the most salient features for QNN inference, thereby surpassing other downsampling approaches. Additionally, we examine state-of-the-art data encoding methods (amplitude, angle, and qubit encoding) and identify amplitude encoding as the optimal strategy, yielding superior accuracy across all tested datasets and qubit configurations. Through computational analyses, we show that our distilled 4-qubit and 8-qubit QNNs achieve competitive performance while utilizing significantly fewer parameters than their classical counterparts. These outcomes highlight the potential of quantum-classical knowledge exchange in enhancing QNN scalability and efficiency. Our results establish a promising paradigm for bridging classical deep learning and emerging quantum computing, paving the way for more powerful, resource-conscious models in quantum machine intelligence.
Keywords Parameterized quantum circuits $\cdot$ Quantum superposition $\cdot$ Knowledge distillation $\cdot$ Neural networks  $\cdot$ Image classification
@@ 1 Introduction
Quantum machine learning (QML) exhibits remarkable potential to transform computational paradigms by leveraging the principles of quantum superposition and entanglement. This can, in specific scenarios, lead to exponential speedups and richer model expressivity than classical counterparts. In particular, quantum neural networks (QNNs) have recently demonstrated compelling advantages, including faster convergence and enhanced classification accuracy, even when constrained to a similar number of trainable parameters as classical networks [1, 2]. Moreover, recent theoretical insights confirm that the effective dimension of well-designed QNNs can exceed that of classical models, hinting at a superior capacity to tackle intricate data distributions [3].
Nonetheless, the nascent status of QML introduces intrinsic obstacles. Modern quantum hardware is restricted by factors such as limited qubit availability, short coherence times, and susceptibility to noise and decoherence [4, 5]. These challenges severely constrain the scale and complexity of QNNs that can be reliably executed, further exacerbated by stochastic outcomes inherent to quantum sampling [6]. As a result, current quantum platforms often demand extensive resources to mitigate errors, leading to a steep computational overhead when attempting to train deeper, more accurate QNNs [7, 8]. A variety of strategies have been devised to alleviate errors and resource limitations in QNNs, including variational error-mitigation schemes, resource-efficient quantum circuit optimizations, and quantum-to-quantum or classical-to-quantum transfer learning methods. Yet, these approaches remain works in progress and face distinct limitations. For instance, neural error mitigation can require large labeled datasets and computational power to train a secondary neural corrector [9], while hybrid variational error mitigation introduces additional parameters that may escalate optimization complexities. Circuit simplifications and feature-map optimizations [10] are often hardware- or application-specific, restricting their broader applicability. Quantum-to-quantum transfer learning presupposes available large-scale quantum models to serve as teachers, whereas classical-to-quantum transfer still primarily depends on classical representations that may not capture complex quantum correlations or entanglement.
Figure 1: Existing approaches of classical to classical and quantum to quantum knowledge distillation, compared to our proposed method: fusion of classical and quantum approaches through classical to quantum knowledge distillation
Knowledge distillation [11] provides a powerful mechanism to sidestep these issues by allowing a large, potentially resource-intensive “teacher” model to impart its learned representations to a smaller “student” model. This paradigm obviates the need for extensive quantum training, as the distilled student network primarily relies on the teacher’s logits to guide its learning. Crucially, distillation can mitigate noise sensitivity, diminish reliance on cumbersome quantum resources, and unveil deeper structural insights in the student network. Consequently, knowledge distillation is emerging as a robust technique for designing and optimizing neural architectures.
While a handful of studies have examined quantum-to-quantum transfer learning or response-based knowledge distillation in QNNs [9], the domain of classical-to-quantum knowledge transfer remains underexplored. Considering the inherent constraints of QML, it is vital to investigate how pre-trained classical models can effectively enhance QNNs through a mere transfer of their outputs. This study thus addresses the following research questions:
@@
  • Research Question 1: Can knowledge be transferred from a classical neural network (teacher) to a quantum neural network (student) through knowledge distillation?
  • Research Question 2: How does this distillation from classical networks impact the performance of a QNN in real-world image classification tasks?
  • Research Question 3: Can such distillation strategies be leveraged to significantly improve resource-constrained quantum models, thereby minimizing the need for large-scale quantum training?
To address these questions, we propose a novel knowledge distillation framework that transfers learned representations from multiple classical teacher models to smaller QNN student models. As illustrated in Fig. 1, this approach bridges the classical and quantum paradigms by utilizing the logits of high-performing convolutional neural networks (CNNs)—LeNet, AlexNet, ResNet, DenseNet, and EfficientNet—to train 4-qubit and 8-qubit variational QNNs. Our experiments span three popular benchmarks: MNIST [12], FashionMNIST [13], and CIFAR10 [14]. We compare quantum student models with and without distillation and further evaluate their performances through thorough statistical significance testing, ablation studies on dimensionality reduction and data encoding strategies, and a comprehensive error analysis. Our principal contributions can be summarized as follows:
  • Hybrid Distillation Framework: We introduce a robust knowledge distillation pipeline that enables the reuse of existing pre-trained classical models for training resource-efficient QNNs, effectively eliminating the need for extensive quantum-specific error mitigation.
  • Systematic Performance Evaluation: We quantify the gains from classical-to-quantum distillation across three datasets and multiple state-of-the-art teacher models, showcasing significant improvements in QNN accuracy.
  • Comparison with State-of-the-art Techniques: Through controlled ablation experiments on dimensionality reduction and data encoding, we identify best practices for efficiently embedding high-dimensional data into few-qubit circuits. Further, our findings are validated by rigorous statistical tests.
  • Generalizability and Future Outlook: We discuss broader implications of the classical-to-quantum transfer paradigm, highlighting how it can empower cross-domain applications and pave the way for more robust quantum-classical hybrid intelligence.
@@ 2 Related Works
2.1 Problems and Challenges in QML
QML aspires to harness the properties of quantum mechanics to revolutionize computation, but the path is restricted with various challenges and obstacles. The concept of quantum supremacy, where quantum systems perform tasks beyond the reach of classical computers, has seen experimental strides as evidenced by [15] and [16]. However, these pioneering steps also underscore the fragility of quantum states and their vulnerability to decoherence, a problem that Shor’s early work [17] sought to address through quantum error correction, a technique still pivotal in contemporary research [18]. The transient stability of qubits, a topic explored by [19], and the intricate challenge of manipulating quantum states, as demonstrated by [20], add layers of complexity to quantum computing.
@@ 2.1 Problems and Challenges in QML
2.2 Error Mitigation and Resource Handling
@@
The pursuit of reliable quantum computing is marked by a need to address the fundamental limits of quantum error mitigation. Works such as those by [28] provide a foundational understanding of the theoretical bounds in quantum systems, crucial for error correction techniques in neural networks. This is further explored through the lens of scalability and statistics of errors in recent advancements within noisy intermediate-scale quantum (NISQ) technologies [29]. Reference-state error mitigation strategies [30] have shown promise in improving the accuracy of quantum computations, which is vital for the fidelity of knowledge distillation processes. Moreover, the extension of computational reach through error mitigation in noisy quantum processors [31] has direct implications for the robust execution of QNNs. The integration of learning-based techniques for quantum error mitigation [32] parallels the adaptive nature of knowledge distillation, suggesting a methodological symmetry that could be harnessed for transferring knowledge. The evaluation and benchmarking of error mitigation techniques are critical for determining the most effective methods for application in QNNs [33]. This selection is further informed by the development of a unified approach to data-driven error mitigation [34], offering a pathway to harnessing empirical data in the error correction process. Techniques for mitigating errors in quantum approximate optimization [35] can also inform the optimization strategies within the knowledge distillation framework.
@@
On the front of resource handling, the importance of optimizing resource efficiencies is underscored by [36], drawing attention to the balance between computational capabilities and resource constraints. This optimization is directly applicable to the development of QNNs, where the allocation and management of quantum resources are crucial. Furthermore, the critical challenges of scaling quantum computation, as outlined by [37], must be addressed to enable the practical implementation of fault-tolerant QNNs. Insights into the cross-discipline theoretical research on quantum computing, such as those presented by [38], highlight the urgency and significance of resource management in the context of knowledge distillation. Additionally, the potential applications of quantum computing in specific domains like renewable energy optimization [39] illustrate the diverse utility and the resource handling strategies that may be required for energy-efficient QNN operations. Differently from previous approaches, we completely nullify the need for QNN optimization using error mitigation or resource handling. Rather, keeping the student quantum model small and static, we depend fully on pre-trained classical models to improve the performance of the QNN, thus addressing the challenges of QML in a new, out-of-the-box method.
2.3 Transfer Learning and Knowledge Transfer in QML
@@
Transfer learning has emerged as a potent strategy in classical machine learning, enabling models trained on one task to be repurposed for another related task [40, 41]. In QML, this approach promises to address scalability and adaptability challenges that are crucial in the practical application of neural-network quantum states (NNQS) [40]. Scalability, a primary concern in QML, is particularly relevant for tasks that require models to extrapolate learned behaviors to larger quantum systems. Protocols inspired by physics have shown that features learned from smaller systems can significantly accelerate and refine the learning process when applied to more complex systems, as evidenced in the one-dimensional transverse field Ising and Heisenberg XXZ models [40]. The versatility of transfer learning is further showcased by its application to the optimization of neural network potentials, exemplified by ANI-1x, which was retrained to near gold-standard quantum mechanical accuracy. This highlights transfer learning’s broad potential across disciplines, including materials science, biology, and chemistry [41, 42].
@@
Integrating classical and quantum neural networks has extended transfer learning protocols into classical-quantum hybrid domains. By combining pre-trained classical networks with variational quantum circuits, researchers have opened up new possibilities for processing high-dimensional data such as images before entangling them within the quantum processing framework [43]. This hybrid approach not only utilizes the strengths of classical networks in data handling but also harnesses the quantum advantages for feature processing, potentially revolutionizing fields like image recognition and quantum state classification [43, 44]. Moreover, the concept of knowledge distillation has been introduced into the realm of QML, allowing the creation of new QNNs through approximate synthesis [9]. This methodology enables the reduction of circuit layers or the alteration of gate sets without the necessity of training from scratch. Empirical analyses suggest that significant reductions in circuit complexity can be achieved while simultaneously improving accuracy, even under noisy conditions, underscoring the effectiveness of transfer learning techniques in the development of QNNs [9]. More recently, hybrid-classical quantum convolutional neural networks have seen applications in brain tumor classification [45], brain lesion detection [46] and pneumonia detection [47].
@@
Quantum Convolutional Neural Networks (QCNNs) are gaining popularity in both quantum and classical data classification recently [53]. The adoption of transfer learning in QML is proposed to circumvent the scalability constraints of quantum circuits in the noisy intermediate-scale quantum (NISQ) era. For instance, research has shown that small QCNNs can effectively solve complex classification tasks by leveraging pre-trained classical CNNs, thus bypassing the need for expansive quantum circuitry [52]. This approach is validated through numerical simulations of QCNNs, which demonstrate a superior classification accuracy over classical models when pre-trained on different datasets. The synergy between classical and quantum networks is evident when a classical CNN, trained on FashionMNIST data, is utilized to enhance the performance of a QCNN for MNIST data classification [53]. Moreover, very recent studies have also calibrated large language models in QNNs [54] and accomplished hybrid classical-quantum transfer learning for text classification [55]. However, differently from [52] and [53], we use knowledge distillation and show empirically that knowledge transfer is possible across homogeneous datasets in quantum models by simply trying to mimic the soft targets of classical CNNs instead of hectic training by freezing the pre-trained model and transferring the knowledge to a QNN during training in the previous approaches.
3 Preliminaries